### v1

@@ Appendix A Additional Tables and Figures
Note: This figure shows parameter estimates from separate regressions of the share of each ad attribute on overall employer values ($V_{j}$), capturing the change in the fraction of ads with each job attribute associated with a one standard deviation increase in employer value. Estimation are done at the employer-cluster level for the five-year period from 2015 to 2019, and are weighted by the number of worker-years in each cluster. We control for occupation and industry using two-digit occupation and two-digit industry fixed effects and location based on deciles of the number of workers in the municipality of the firm.
@@
Note: This figure documents adjusted $R^{2}$s from regressions of estimated overall ($V_{j}$), pay ($\psi_{j}$) and non-pay ($a_{j}$) employer values on pay and non-pay job attributes and other ad characteristics detected in the text analysis as in Figure 6, decomposing the overall $R^{2}$s into ten broad categories of pay and non-pay attributes. Estimations are done at the employer-cluster level for the five-year period from 2015 to 2019, and are weighted by the number of worker-years in each cluster.
@@
Note: This figure documents adjusted $R^{2}$ from regressions of estimated population group-specific model parameters on ad attributes and characteristics in the text analysis as in Figure 7. Estimations are done at the employer-cluster level for each of the four five-year period from 2000 to 2019, and are weighted by the total number of worker-years in each cluster and period. Overall $R^{2}$s are from regressions that include ad attributes and other ad characteristics, while partial $R^{2}$s are from regressions that also control for the composition of industry, occupation and location in each cluster. Industry and occupation controls are defined at the two digit level. Location indicators split local municipalities into deciles, with municipalities in the same group having similar number of workers (Oslo as an own group).
@@
Note: This figure documents adjusted $R^{2}$ from regressions of estimated population group-specific model parameters on ad attributes and characteristics in the text analysis as in Figure 7, with fixed weights across sample splits. Estimations are done at the employer-cluster level for the five-year period from 2015 to 2019, separately by gender (Panels A-B) and education (Panels C-D), and are weighted by the total number of worker-years in each cluster, as opposed to by cluster and sub-sample as in Figure 7. Overall $R^{2}$s are from regressions that include ad attributes and other ad characteristics, while partial $R^{2}$s are from regressions that also control for the composition of industry, occupation and location in each cluster. Industry and occupation controls are defined at the two digit level. Location indicators split local municipalities into deciles, with municipalities in the same group having similar number of workers (Oslo as an own group).
@@ B.2 Extracting Common Attributes from “We Offer” Lists
After estimating the model, we label the topics into four identified topics by considering the word distributions and the text (list) specific topic mixtures. Table B.1 documents the ten words with the highest association with each topic, i.e., the highest probability of being drawn in each topic. Words in topics 1-3 are associated with 1) (both pay and non-pay) job amenities, 2) skill requirements, and 3) task descriptions, while words in topic 4 contains a mixture of requirements (e.g. education and police certificate), additional information (e.g. contact information), and ambiguous words (numbers and “work”). We label the topics 1-3 according to these observations, and topic 4 as a “Residual Category”.
@@
We validate these labels by inspecting the lists with the highest association with each topic in Figure B.1-B.4. These lists serve as stylized examples of each topic and omit only considering one word at the time. These lists largely confirms the labels associated with each topic. The list most strongly associated with the “We Offe”-topic starts with the header “we offer:” and contains information about various non-pay amenities. The list most strongly associated with the ”Residual Category”-topic contains various information about the job, such as the hiring process and and information about pay. Overall, these lists suggests that lists in the “We Offer” category contain dense information about amenities, although we are likely to lose some information about pay and non-pay in lists sorted into other categories (e.g., information about pay in lists assigned to the “Residual Category”).
@@
Next, we extract and inspect the 200 most common unigrams, bigrams and trigrams, excluding phrases that includes stopwords (e.g., “and”). We add phrases that indicate (pay or non-pay) job amenities to our dictionary of phrases, and create new attribute categories whenever a phrase indicates an attribute not currently in our list.
@@ B.3 Description of Detailed Pay and Non-Pay attributes
Tables B.3 to B.5 provides descriptions of each of the attribute we search for in job ads, as well as examples of phrases we use to detect these attributes.
@@ Appendix D Predictive Power of Publicly Posted Job Attributes
Which job attributes contain the most predictive power? In Section 5.2 we decompose the overall $R^{2}$ from regressions of employer values on ad attributes by sequentially adding groups of attributes. This approach abstracts away from the 47 individual attributes and four other ad characteristics, and depends on the order of which attributes are added. This section documents that the main patterns in Section 5.2 and Appendix Figure A.6 hold when we relax the dependency on the ordering and focus on individual attributes and ad characteristics.
@@
For regressions without additional controls, we find that attributes related to convenient or inconvenient hours are important predictors (i.e., among the first five added variables) for all model components. This confirms the results in Section 5.2. For pay values, we find that attributes related to financial attributes and hours of work are also important. For non-pay values, we find that “on-the-job training” and “permanent job” are also among the most important predictors, together with “variation in tasks” and the number of words in ads. This confirms the observation that non-pay values are related to different types of attributes posted in job ads (Section 5.2), and uncovers that “shift work” carries a significant portion of the prediction power among inconvenient hours attributes. The pattern for job security is less clear, with no real prediction value from attributes related to contract duration.
For regressions with additional controls, we find a somewhat different pattern. Reassuringly, “permanent job” is the first attribute added in regressions of job security with additional controls. Yet, again, most of the predictive power stems from attributes related to convenient or inconvenient hours, besides financial attributes and hours of work.
@@
No Additional Controls (Adjusted $R^{2}$) With Additional Controls (Partial $R^{2}$)
@@
– Competitive Pay ∗∗ ∗∗
– Collective Agreement Pay ∗∗
@@
– Pension Scheme ∗∗∗ ∗∗∗
– Insurance Scheme ∗∗∗ ∗∗∗ ∗∗∗
@@
– Good Career Paths ∗∗
– On-the-Job Training ∗∗∗ ∗∗∗
@@
– Part-time Contract ∗∗ ∗∗∗ ∗∗∗ ∗∗∗ ∗∗∗
– Full-time/Part-time Choice ∗∗∗ ∗∗ ∗∗ ∗∗∗ ∗∗
– On-call Employment
@@
– Regular Daytime Work Schedule ∗∗∗ ∗∗∗ ∗∗∗ ∗∗∗ ∗∗∗ ∗∗∗
– Exempt from Work Hour Regulations ∗∗∗ ∗∗∗
@@
– Shift Work ∗∗∗ ∗∗∗ ∗∗∗ ∗∗∗ ∗∗ ∗∗
– Weekend/Evening/Night Work ∗∗∗ ∗∗ ∗∗
@@
– Possibility for Remote Work ∗∗ ∗∗
– Shared Office Space
@@
– Challenging Tasks
– Variation in Tasks ∗∗∗ ∗∗ ∗∗∗
– Responsibility in Job ∗∗
– Independence in Performing Tasks ∗∗ ∗∗
@@
– Parking Space On Premises
– Company Vehicle ∗∗
– Any Welfare Scheme ∗∗ ∗∗∗ ∗∗∗
– Company Cabin
@@
– Number of Words ∗∗ ∗∗∗ ∗∗∗
– Posted Online ∗∗ ∗∗
@@
– Number of Attributes ∗∗
∗∗∗ Model with 5 regressors, ∗∗ Model with 10 regressors, Model with 20 regressors.

### v2

@@ Appendix A Additional Tables and Figures
Note: This figure shows parameter estimates and $90\%$ confidence intervals from separate regressions of the share of each ad attribute on overall employer values ($V_{j}$), capturing the change in the fraction of ads with each job attribute associated with a one standard deviation increase in employer value. Estimation are done at the employer-cluster level for the five-year period from 2015 to 2019, and are weighted by the number of worker -years in each cluster. We control for occupation and industry using two-digit occupation and two-digit industry fixed effects and location based on deciles of the number of workers in the municipality of the firm.
Figure A.6: Publicly Advertised Job Attributes By Employer Values: Smaller Clusters.
(a) Pay ($\psi_{j}$)
(b) Non-Pay ($a_{j}$)
Note: This figure shows parameter estimates and $90\%$ confidence intervals from separate regressions of the share of each ad attribute on employer pay ($\psi_{j}$) and non-pay ($a_{j}$) values, capturing the change in the fraction of ads with each job attribute associated with a standard deviation increase in employer value. Estimations are done at the employer-cluster level for the five-year period from 2015 to 2019, and are weighted by the number of worker-years. We control for occupation and industry using two-digit occupation and two-digit industry fixed effects and location based on deciles of the number of workers in the municipality of the firm (Oslo as an own group). On average, each cluster has 25 establishments, i.e., $G=$round$(J/25)$, while Figure 5 shows the corresponding results in our baseline with $G=$round$(J/50)$.
Figure A.7: Publicly Advertised Job Attributes By Employer Values: Larger Clusters.
(a) Pay ($\psi_{j}$)
(b) Non-Pay ($a_{j}$)
Note: This figure shows parameter estimates and $90\%$ confidence intervals from separate regressions of the share of each ad attribute on employer pay ($\psi_{j}$) and non-pay ($a_{j}$) values, capturing the change in the fraction of ads with each job attribute associated with a standard deviation increase in employer value. Estimations are done at the employer-cluster level for the five-year period from 2015 to 2019, and are weighted by the number of worker-years. We control for occupation and industry using two-digit occupation and two-digit industry fixed effects and location based on deciles of the number of workers in the municipality of the firm (Oslo as an own group). On average, each cluster has 100 establishments, i.e., $G=$round$(J/100)$, while Figure 5 shows the corresponding results in our baseline with $G=$round$(J/50)$.
@@
Note: This figure documents adjusted $R^{2}$s from regressions of estimated employer values on pay and non-pay job attributes and other ad characteristics detected in the text analysis as in Figure 6, decomposing the overall $R^{2}$s into ten broad categories of pay and non-pay attributes. Estimations are done at the employer-cluster level for the five-year period from 2015 to 2019, and are weighted by the number of worker-years in each cluster.
Figure A.9: Predictive Power of Publicly Advertised Job Attributes: By Industry.
Note: This figure documents adjusted $R^{2}$s from regressions of estimated overall employer values on pay and non-pay job attributes and other ad characteristics detected in the text analysis as in Figure 6. Estimations are done separately for the ten main industry groups at the employer-cluster level for the five-year period from 2015 to 2019, and are weighted by the number of worker-years in each cluster. We assign an industry to each cluster $g$ based on the modal industry among the establishments in the cluster.
Figure A.10: Predictive Power of Publicly Advertised Job Attributes: By Cluster Size.
(a) Baseline Model: $G=\text{round}(J/50)$
(b) Smaller Clusters: $G=\text{round}(J/25)$
(c) Larger Clusters: $G=\text{round}(J/100)$
Note: This figure documents adjusted $R^{2}$ from regressions of cluster-size specific estimated model parameters on ad attributes and characteristics in the text analysis as in Figure 6. Estimations are done at the employer-cluster level for each of the four five-year period from 2000 to 2019, and are weighted by the total number of worker-years in each cluster and period. Overall $R^{2}$s are from regressions that include ad attributes and other ad characteristics, while partial $R^{2}$s are from regressions that also control for the composition of industry, occupation and location in each cluster. Industry and occupation controls are defined at the two-digit level. Location indicators split local municipalities into deciles, with municipalities in the same group having similar number of workers (Oslo as an own group).
@@
Note: This figure documents adjusted $R^{2}$ from regressions of estimated population group-specific model parameters on ad attributes and characteristics in the text analysis as in Figure 6. Estimations are done at the employer-cluster level for each of the four five-year period from 2000 to 2019, and are weighted by the total number of worker-years in each cluster and period. Overall $R^{2}$s are from regressions that include ad attributes and other ad characteristics, while partial $R^{2}$s are from regressions that also control for the composition of industry, occupation and location in each cluster. Industry and occupation controls are defined at the two digit level. Location indicators split local municipalities into deciles, with municipalities in the same group having similar number of workers (Oslo as an own group). The sample corresponds to the last rows in Panels A-D in Appendix Table A.1
@@
Note: This figure documents adjusted $R^{2}$ from regressions of estimated population group-specific model parameters on ad attributes and characteristics in the text analysis as in Figure 7, with fixed weights across sample splits. Estimations are done at the employer-cluster level for the five-year period from 2015 to 2019, separately by gender (Panels A-B) and education (Panels C-D), and are weighted by the total number of worker-years in each cluster, as opposed to by cluster and population group as in Figure 7. Overall $R^{2}$s are from regressions that include ad attributes and other ad characteristics, while partial $R^{2}$s are from regressions that also control for the composition of industry, occupation and location in each cluster. Industry and occupation controls are defined at the two digit level. Location indicators split local municipalities into deciles, with municipalities in the same group having similar number of workers (Oslo as an own group).
@@ B.2 Extracting Common Attributes from “We Offer” Lists
After estimating the model, we label the topics into four identified topics by considering the word distributions and the text (list) specific topic mixtures. Table B.1 documents the ten words with the highest association with each topic, i.e., the highest probability of being drawn in each topic. Words in topics 1-3 are associated with 1) pay and non-pay job attributes (amenities), 2) skill requirements, and 3) task descriptions, while words in topic 4 contains a mixture of requirements (e.g. education and police certificate), additional information (e.g. contact information), and ambiguous words (numbers and “work”). We label the topics 1-3 according to these observations, and topic 4 as a “Residual Category”.
@@
We validate these labels by inspecting the lists with the highest association with each topic in Figure B.1-B.4. These lists serve as stylized examples of each topic and omit only considering one word at the time. These lists largely confirms the labels associated with each topic. The list most strongly associated with the “We Offer”-topic starts with the header “we offer:” and contains information about various non-pay amenities. The list most strongly associated with the “Residual Category”-topic contains various information about the job, such as the hiring process and and information about pay. Overall, these lists suggests that lists in the “We Offer” category contain dense information about amenities, although we are likely to lose some information about pay and non-pay in lists sorted into other categories (e.g., information about pay in lists assigned to the “Residual Category”).
@@
Next, we extract and inspect the 200 most common unigrams, bigrams and trigrams, excluding phrases that includes stopwords (e.g., “and”). We add phrases that indicate pay or non-pay job attributes (amenities) to our dictionary of phrases, and create new attribute categories whenever a phrase indicates an attribute not currently in our list.
@@ B.3 Description of Detailed Pay and Non-Pay Attributes
Tables B.3 to B.5 provides descriptions of each of the attribute we search for in job ads, as well as examples of phrases we use to detect these attributes. In our final dataset of job ads, we label an attribute as detected using an indicator variable whenever the text of the associated job ad contains any of the phrases associated with this attribute. [^45] The exception to this rule is the attribute “Any Other Mention of Pay”, which we construct by first searching for a set of generic words associated with pay and labelling the attribute as detected in the associated job ad, but later on set the indicator for whether this attribute is detected to zero if any of the other attributes listed under “Compensation Scheme” were already detected by the same set of word.
@@ Appendix D Predictive Power of Publicly Posted Job Attributes
Which job attributes hold the highest predictive power about the actual employer pay and non-pay values? In Section 5.2, we decompose the overall $R^{2}$ from regressions of employer values on ad attributes by sequentially adding groups of attributes. This approach abstracts away from the 47 individual attributes and four other ad characteristics, and the results of this approach may thus depend on the grouping and ordering of attributes. This section documents that the main patterns in Section 5.2 and Appendix Figure A.8 hold when we relax the dependency on the ordering and focus on individual attributes and ad characteristics.
@@
For regressions without additional controls, we find that attributes related to convenient or inconvenient hours are important predictors (i.e., among the first five added variables) for all model components. This confirms the results in Section 5.2. For pay values, we find that attributes related to financial attributes and hours of work are important, but also “regular daytime work schedule” and “shift work”. For non-pay values, we also find that “regular daytime work schedule”, “shift work” and “weekend/evening/night work” are important, but also “social environment” and “permanent job”. This confirms the observation that non-pay values are related to particular types of attributes posted in job ads (Section 5.2), and uncovers that “regular daytime work schedule” and “shift work” carry a significant portion of the prediction power among convenient and inconvenient hours attributes. Finally, for job security, the presence of “any welfare scheme” also appears to be important, which may signal that the company has in place specific schemes to value their workers’ welfare.
For regressions with additional controls, we find broadly similar patterns. The various attributes related to convenient and inconvenient hours again feature prominently among those with the highest predictive power, while “permanent job” is now reassuringly the first attribute added in regressions for job security with the additional controls.
@@
[…]
@@
– Competitive Pay
– Collective Agreement Pay
@@
– Pension Scheme ∗∗ ∗∗∗
– Insurance Scheme ∗∗∗ ∗∗∗ ∗∗∗ ∗∗∗
@@
– Good Career Paths ∗∗ ∗∗
– On-the-Job Training
@@
– Part-time Contract ∗∗∗ ∗∗ ∗∗ ∗∗∗
– Full-time/Part-time Choice ∗∗∗ ∗∗ ∗∗ ∗∗
@@
– Regular Daytime Work Schedule ∗∗∗ ∗∗∗ ∗∗∗ ∗∗∗ ∗∗∗ ∗∗∗ ∗∗∗ ∗∗∗
– Exempt from Work Hour Regulations ∗∗∗ ∗∗∗ ∗∗
@@
– Shift Work ∗∗ ∗∗∗ ∗∗∗ ∗∗∗
– Weekend/Evening/Night Work ∗∗∗ ∗∗∗ ∗∗∗ ∗∗∗
– On-call Employment
@@
– Possibility for Remote Work ∗∗
– Shared Office Space ∗∗
@@
– Challenging Tasks ∗∗
– Variation in Tasks ∗∗ ∗∗
– Responsibility in Job
– Independence in Performing Tasks
@@
– Parking Space On Premises ∗∗ ∗∗
– Company Vehicle
– Any Welfare Scheme ∗∗ ∗∗∗ ∗∗ ∗∗∗
– Company Cabin ∗∗
@@
– Number of Words ∗∗ ∗∗ ∗∗∗ ∗∗∗
– Posted Online ∗∗
@@
– Number of Attributes
∗∗∗ Model with 5 regressors, ∗∗ Model with 10 regressors, model with 20 regressors